OpenStudy (anonymous):

3x2 + 17x + 10 how would I factor this completely?

OpenStudy (anonymous):

(3x+2)(x+5)

OpenStudy (anonymous):

3*10=30 2+15=17 2*15=30 write 17x=2x+15x and then make factors.

OpenStudy (anonymous):

using the a-c method so you find factors of a*c that add up to b: \[3x^2+15x+2x+10\] now factor by grouping so: \[3x(x+5)+2(x+5)\] \[(3x+2)(x+5)\]